# Recurrent convolutional neural networks for non-adiabatic dynamics of quantum-classical systems

### v1

@@ Abstract
Recurrent neural networks (RNNs) have recently been extensively applied to model the time-evolution in fluid dynamics, weather predictions, and even chaotic systems thanks to their ability to capture temporal dependencies and sequential patterns in data. Here we present a RNN model based on convolutional neural networks for modeling the nonlinear non-adiabatic dynamics of hybrid quantum-classical systems. The dynamical evolution of the hybrid systems is governed by equations of motion for classical degrees of freedom and von Neumann equation for electrons. The physics-aware recurrent convolutional (PARC) neural network structure incorporates a differentiator-integrator architecture that inductively models the spatiotemporal dynamics of generic physical systems. Validation studies show that the trained PARC model could reproduce the space-time evolution of a one-dimensional semi-classical Holstein model with comparable accuracy to direct numerical simulations. We also investigate the scaling of prediction errors with size of training dataset, prediction window, step-size, and model size.
@@ I Introduction
Recurrent Neural Networks (RNNs) are an established framework for analyzing sequential or dynamic data with applications across physics [1, 2, 3, 4]. Unlike feed-forward networks, which assume independence between inputs, RNNs incorporate internal feedback loops which can capture dependencies between sequential data. This is a crucial behavior when modeling the evolution of time-dependent physical systems and trajectories with strong temporal correlations between data points. RNNs can learn the underlying dynamics from time-series data without requiring explicit knowledge of governing equations. They can predict future states of a system based on its past behavior. Indeed, RNN is among one of the prominent deep-learning approaches often employed to solve partial differential equations (PDEs) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], which are the ubiquitous tools for modeling the spatiotemporal evolution of most physical systems.
@@
The essential idea behind the RNN approach to PDE, namely using current and previous step states to calculate the state at the next time step, is actually similar to most time-stepping numerical methods for solving time-dependent PDEs, such as Euler’s, Crank-Nicolson, and higher-oder Runge-Kutta schemes [19, 20, 21]. The RNN serves as a surrogate time-stepping method replacing the numerical finite-difference or finite-element schemes. The numerical solution on each of the grid point (for finite difference) or grid cell (for finite element) computed at a set of contiguous time points can be treated as a time sequence data for training a multi-input RNN. Standard supervised learning methods are used to optimize the parameters of neural network layers through learning the numerical solutions on the grids.
@@
The dynamics of physical systems is also characterized by the emergence of complex spatial patterns, such as stripes and vortices. This indicates the importance of spatiotemporal correlations in the dynamical modeling of such physical systems. Yet, traditional neural networks often perform poorly at capturing local relationships due to their fully-connected nature, Convolutional Neural Networks (CNNs) [22], on the other hand, excel at such tasks. CNNs are a variety of neural network which utilize convolution layers to process a grid or array of values using a ”sliding” kernel to extract learned features with geometric structure. Convolution layers strongly exploit patterns with geometric locality in inputs, and within a CNN many convolution layers are often composed to extract a hierarchy of features starting from low-level patterns to high-level and complex relationships. Because of these unique capabilities, CNNs have become a crucial component in deep-learning approach to solving PDEs [15, 16, 17, 18].
@@
The recently proposed RNN, dubbed physics-aware recurrent convolution (PARC) [17, 18], further incorporates physics information into its architecture, thus offering improved accuracy and representation capability. Indeed, machine learning (ML) models with physical constrained explicitly incorporated have attracted great interest among researchers. For example, the ML force-field framework for ab initio molecular dynamics utilizes the locality principle to implement local energy or force models, enabling linear scaling computation [23, 24, 25, 26, 27, 28, 29]. Another example is the equivariant neural networks which allows physical symmetries to be explicitly included on a supervised learning model. Their architecture is designed in such a way that the neurons of each layer exhibit well-defined transformation properties under operation of a given symmetry group [30, 31, 32, 33].
@@
In this work we propose a deep-learning approach based on the PARC architecture for modeling the non-adiabatic dynamics of lattice models with coexisting classical and quantum electron degrees of freedom. Such hybrid quantum-classical models are an important simplification for modeling realistic physical systems with, e.g. multiple length and time scales. One of the most prominent examples is the ab initio molecular dynamics methods widely used in quantum chemistry and materials science [34]. Yet, direct dynamical simulation of such hybrid systems is still a challenging computational task because of the exponentially large dimension of the Hilbert space associated with the quantum subsystem.
Even in the absence of direct electron-electron interactions, a many-body description of electrons is still required in order to accommodate the Fermi-Dirac quantum statistics. In the non-interacting case, the complexity of simulating a many-electron wave function can be reduced to the dynamics of single-electron density matrix or correlation functions $\rho(\mathbf{r},\mathbf{r}^{\prime},t)$ where $\mathbf{r}$, $\mathbf{r}^{\prime}$ are coordinates of lattice sites. As a result, for a $D$-dimensional system $\mathbf{r}\in\mathbb{Z}^{D}$, the minimum description in terms of density matrix requires a $2D$-dimensional time-dependent differential equations. As a proof-of-principle, here we consider the 1-dimensional (1D) semiclassical Holstein model and build a PARC framework to model its non-adiabatic dynamics under a quantum quench. By employing a recurrent next-step prediction approach, we show that the PARC model can accurately perform multi-step integration from any given starting point. Furthermore, we examine error accumulation during successive recurrent predictions, as well as the scaling effects of data step size, training set size, and model size on prediction accuracy.
Although the dynamics of discrete lattice systems is formally described by coupled ordinary differential equations (ODEs), spatial couplings between dynamical variables nonetheless resemble those of discretized PDEs using either finite-difference or finite-element schemes. The spatial correlations and potential emergent spatial patterns can be efficiently captured by the CNN built in PARC, as discussed above. Indeed, similar CNN-based force-field models have been developed for spin dynamics of the so-called s-d system, a lattice model of metallic magnets [35]. Importantly, the convolution operation with a finite-sized kernel naturally incorporates the locality principle into the PARC model, which can be straightforwardly scaled to larger systems.
@@
The rest of the paper is organized as follows. In Sec. II, we discuss the non-adiabatic dynamics of the semiclassical Holstein model, within the Ehrenfest dynamics framework. After a brief review on the PARC formulation, the modified recurrent architecture for modeling the dynamics of Holstein model is presented in Sec. III. The results and several benchmark analyses are discussed in Sec. IV. Finally we present a summary and outlook in Sec. V.
@@ II Non-adiabatic dynamics of Holstein model
^$\hat{\mathscr{H}}=-t_{\rm nn}\sum_{i}\left(\hat{c}^{{\dagger}}_{i}\hat{c}^{\,}_{i+1}+\hat{c}^{{\dagger}}_{i+1}\hat{c}^{\,}_{i}\right)-g\sum_{i}\left(\hat{n}_{i}-\frac{1}{2}\right)\hat{Q}_{i}$
@@
Here $\hat{c}^{\dagger}_{i}$ ($\hat{c}^{\,}_{i}$) denotes the creation (annihilation) operator of an electron at lattice site-$i$, $Q_{i}$ represents a local phonon degree of freedom, $P_{i}$ is the associated conjugate momentum, $m$ is an effective mass, $\Omega$ is the intrinsic oscillation frequency, and $K\equiv m\Omega^{2}$ is the force constant, finally $g$ denotes the electron-phonon coupling coefficient. The first term ^$\hat{\mathscr{H}}_{\mathrm{e}}$ describes the hopping of electrons between nearest neighboring sites. The second part essentially describes the Einstein phonon model, which corresponds to a set of simple harmonic oscillators each associated with a lattice site. The third term describes a local coupling between the electron density $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}^{\,}_{i}$ and the displacement of the oscillator.
The Holstein model at half-filling on various bipartite lattices exhibits a robust commensurate CDW order that breaks the sublattice symmetry [39, 40, 41, 42]. In one dimension, this commensurate CDW order is characterized by a ultra-short period modulation of electron density,
$\langle n_{i}\rangle=\overline{n}+\delta n\cos\left(\frac{\pi}{a}x_{i}\right),$ (2)
@@
Here we employ the Ehrenfest dynamics framework [43, 34] to describe the semiclassical dynamics of the Holstein model. A similar semiclassical dynamics method was recently employed to study the photo-emission and long-time behaviors of CDW states in the 1D Holstein model [44]. To this end, we assume a product form for the quantum state of the system: |$\ket{\Gamma(t)}=\ket{\Phi(t)}\otimes\ket{\Psi(t)}$, where |$\ket{\Phi(t)}$ and |$\ket{\Psi(t)}$ denote the phonon and electron wave-functions, respectively. The semi-classical approximation for the lattice subsystem amounts to a direct product wave function |$\ket{\Phi(t)}=\prod_{i}\ket{\phi_{i}(t)}$ for the phonons. As a result, the expectation value of phonon operators, e.g. $\langle\Gamma(t)|\hat{Q}_{i}|\Gamma(t)\rangle$ reduces to $Q_{i}(t)\equiv\langle\phi_{i}(t)|\hat{Q}_{i}|\phi_{i}(t)\rangle$, and similarly for the momentum operators, $P_{i}(t)\equiv\langle\Gamma(t)|\hat{P}_{i}|\Gamma(t)\rangle=\langle\phi_{i}(t)|\hat{P}_{i}|\phi_{i}(t)\rangle$.
The dynamics of these ‘classical’ variables is given by the expectation of Heisenberg equation, e.g. $d\langle\hat{P}_{i}\rangle/dt=-i\langle[\hat{P}_{i},\hat{\mathscr{H}}]\rangle/\hbar$, where delimited-⟨$\langle$..$\rangle$ … ⟩ is the expectation value computed using the full wave-function |$\ket{\Gamma(t)}$. Direct calculation of the commutators yields the coupled Hamiltonian dynamics
@@
There are two characteristic time scales for the dynamics of the Holstein model. First, from the bandwidth of the electron tight-binding model $W=4t_{\rm nn}$, one can define a time scale $\tau_{\rm e}=\hbar/t_{\rm nn}$ for the electron dynamics. Another time scale is given by the natural frequency $\Omega$ of the local simple harmonic oscillator: $\tau_{\rm L}=1/\Omega$. The dimensionless adiabatic parameter is defined as the ratio $r=\tau_{\rm e}/\tau_{\rm L}=\hbar\Omega/t_{\rm nn}$. The electron-phonon coupling can also be characterized by a dimensionless parameter $\lambda$ as follows. First, let $Q^{*}$ be lattice distortions estimated from the balance of elastic energy and electron-phonon coupling: $KQ^{*2}\sim g\langle n\rangle Q^{*}$. Assuming electron number $\langle n\rangle\sim 1$, we obtain $Q^{*}\sim g/K$. The dimensionless parameter $\lambda=gQ^{*}/W=g^{2}/WK$ corresponds to the ratio of electron-phonon coupling to the bandwidth. For all simulations discussed below, the adiabatic parameter is set to $r=0.4$. The simulation time is measured in unit of $\tau_{\rm e}$, energies are measured in units of $t_{\rm nn}$, and the lattice distortion is expressed in terms of $Q^{*}$.
III PARC neural network for Holstein Model
In this section we present a PARC neural network for modeling the non-adiabatic dynamics of the 1D semiclassical Holstein model. While the state of the simple harmonic oscillators on a chain is described by two 1D arrays $\{Q_{i}\}$ and $\{P_{i}\}$, the minimum dynamical variables for the many-electron system are given by a density matrix $\rho_{ij}$, due to the quantum Fermi-Dirac statistics. This means that, for the simpler case of non-interacting electrons, the 1D quantum dynamics can be effectively modeled by an effective 2D classical dynamics.
@@ III PARC neural network for Holstein Model
Figure 1: Schematic diagram of the recurrent structure for non-adiabatic dynamics of the semiclassical Holstein model. The framework is based on three PARC neural networks, one for each of the three dynamical degrees of freedom: $Q$, $P$, and $\rho$. Each of the three PARC neural nets is composed of a differentiator CNN and a integrator CNN.
III.1 PARC: a brief review
We first briefly review the basic principles of PARC, our presentation closely follows that in Ref. [17]. By introducing a state vector $\mathbf{u}=[Q_{i},P_{i},\rho_{ij}]$, the semiclassical dynamics of the Holstein model can be cast into a form of the standard ordinary differential equation:

### v3

# Recurrent convolutional neural networks for modeling non-adiabatic dynamics of quantum-classical systems

@@ Abstract
Recurrent neural networks (RNNs) have recently been extensively applied to model the time-evolution in fluid dynamics, weather predictions, and even chaotic systems thanks to their ability to capture temporal dependencies and sequential patterns in data. Here we present a RNN model based on convolutional neural networks for modeling the nonlinear non-adiabatic dynamics of hybrid quantum-classical systems. The dynamical evolution of the hybrid systems is governed by equations of motion for classical degrees of freedom and von Neumann equation for electrons. The physics-aware recurrent convolution (PARC) neural network structure incorporates a differentiator-integrator architecture that inductively models the spatiotemporal dynamics of generic physical systems. We apply our RNN approach to learn the space-time evolution of a one-dimensional semi-classical Holstein model after an interaction quench. For shallow quenches (small changes in electron-lattice coupling), the deterministic dynamics can be accurately captured using a single-CNN-based recurrent network. In contrast, deep quenches induce chaotic evolution, making long-term trajectory prediction significantly more challenging. Nonetheless, we demonstrate that the PARC-CNN architecture can effectively learn the statistical climate of the Holstein model under deep-quench conditions.
@@ I Introduction
Recurrent Neural Networks (RNNs) are an established framework for analyzing sequential or dynamic data with applications across physics [1, 2, 3, 4]. Unlike feed-forward networks, which assume independence between inputs, RNNs incorporate internal feedback loops which can capture dependencies between sequential data. This is a crucial behavior when modeling the evolution of time-dependent physical systems and trajectories with strong temporal correlations between data points. RNNs can learn the underlying dynamics from time-series data without requiring explicit knowledge of governing equations. They can predict future states of a system based on its past behavior. Indeed, RNNs are among one of the prominent deep-learning approaches often employed to solve partial differential equations (PDEs) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], which are the ubiquitous tools for modeling the spatiotemporal evolution of most physical systems.
@@
The essential idea behind the RNN approach to PDE, namely using current and previous step states to calculate the state at the next time step, is actually similar to most time-stepping numerical methods for solving time-dependent PDEs, such as Euler’s, Crank-Nicolson, and higher-order Runge-Kutta schemes [19, 20, 21]. The RNN serves as a surrogate time-stepping method replacing the numerical finite-difference or finite-element schemes. The numerical solution on each of the grid point (for finite difference) or grid cell (for finite element) computed at a set of contiguous time points can be treated as neural network input in the form of one time sequence of data. The deep learning framework is then trained to predict any time-dependent PDEs from the time series data supported on the grids. In other words, the supervised training process can be viewed as a way for the deep learning framework to learn the numerical solution from the input data by optimizing the coefficients of the neural network layers.
A central theme in recent research - particularly motivated by applications to chaotic dynamics in many physical systems - is learning the long-term climate of chaotic systems, characterized by ergodic statistics such as attractor geometry and Lyapunov spectra, rather than insisting on accurate long-horizon trajectories for individual initial conditions [22, 23]. For canonical low- and moderate-dimensional benchmarks, as well as spatiotemporal PDEs, echo state networks have been trained to reproduce climate measures - including Lyapunov spectra - for systems such as Lorenz and the Kuramoto-Sivashinsky equation, with scalability to large spatial domains via parallel reservoir architectures [22, 24]. Robustness analyses across multiple training realizations and network topologies reveal substantial variability, with runs achieving better short-term accuracy also tending to reproduce the long-term climate more faithfully [23]. Similarly, in transitional shear flows, models trained exclusively on turbulent data can nevertheless capture laminar behavior and the statistics of turbulent-laminar switching [25]. This emphasis on statistical fidelity spans diverse implementations, from large parallel reservoirs to minimal physical reservoirs - such as a single driven pendulum - capable of performing both temporal and non-temporal tasks by exploiting transient dynamics [26].
@@
The dynamics of physical systems is also characterized by the emergence of complex spatial patterns, such as stripes and vortices. This indicates the importance of spatiotemporal correlations in the dynamical modeling of such physical systems. Yet, traditional neural networks often perform poorly at capturing local relationships due to their fully-connected nature; Convolutional Neural Networks (CNNs) [27], on the other hand, excel at such tasks. CNNs are a variety of neural network which utilize convolution layers to process a grid or array of values using a “sliding” kernel to extract learned features with geometric structure. Convolutional layers strongly exploit patterns with geometric locality in inputs, and within a CNN many convolutional layers are often composed to extract a hierarchy of features starting from low-level patterns to high-level and complex relationships. Because of these unique capabilities, CNNs have become a crucial component in deep-learning approaches to solving PDEs [15, 16, 17, 18].
@@
A recently proposed RNN scheme, dubbed physics-aware recurrent convolution (PARC) [17, 18], further incorporates physics information into its architecture, thus offering improved accuracy and representation capability. Indeed, machine learning (ML) models with physical constraints explicitly incorporated have attracted great interest among researchers. For example, the ML force-field framework for ab initio molecular dynamics utilizes the locality principle to implement local energy or force models, enabling linear scaling computation [28, 29, 30, 31, 32, 33, 34]. Another example is the equivariant neural networks which allows physical symmetries to be explicitly included on a supervised learning model. Their architecture is designed in such a way that the neurons of each layer exhibit well-defined transformation properties under operation of a given symmetry group [35, 36, 37, 38].
@@
In this work we propose a deep-learning approach based on both a standard CNN and the PARC architecture for modeling the non-adiabatic dynamics of lattice models with coexisting classical and quantum electron degrees of freedom. Such hybrid quantum-classical models are an important simplification for modeling realistic physical systems with, e.g. multiple length and time scales. One of the most prominent examples is the ab initio molecular dynamics methods widely used in quantum chemistry and materials science [39]. Yet, direct dynamical simulation of such hybrid systems is still a challenging computational task because of the exponentially large dimension of the Hilbert space associated with the quantum subsystem.
Even in the absence of direct electron-electron interactions, a many-body description of electrons is still required in order to accommodate the Fermi-Dirac quantum statistics. In the non-interacting case, the complexity of simulating a many-electron wave function can be reduced to the dynamics of single-electron density matrix or correlation functions $\rho(\mathbf{r},\mathbf{r}^{\prime},t)$ where $\mathbf{r}$, $\mathbf{r}^{\prime}$ are coordinates of lattice sites. As a result, for a $D$-dimensional system $\mathbf{r}\in\mathbb{Z}^{D}$, the minimum description in terms of density matrix requires $D^{2}$-dimensional time-dependent differential equations. As a proof-of-principle, here we consider the 1-dimensional (1D) semi-classical Holstein model and demonstrate a CNN-based framework for its non-adiabatic dynamics under a quantum quench.
Although the dynamics of discrete lattice systems is formally described by coupled ordinary differential equations (ODEs), spatial couplings between dynamical variables nonetheless resemble those of discretized PDEs using either finite-difference or finite-element schemes. The spatial correlations and potential emergent spatial patterns can be efficiently captured by a CNN, as discussed above. Indeed, similar CNN-based force-field models have been developed for spin dynamics of the so-called s-d system, a lattice model of metallic magnets [40]. Importantly, the convolution operation with a finite-sized kernel naturally incorporates the locality principle, which can be straightforwardly scaled to larger systems.
@@
The remainder of the paper is organized as follows. Section II introduces the non-adiabatic dynamics of the semi-classical Holstein model within the Ehrenfest dynamics framework. In Section III, we describe the formulation, implementation, and results of a standard CNN model applied to the Holstein model under a shallow quench, demonstrating that even a compact CNN can accurately capture the time evolution. Section IV presents the formulation and implementation of a PARC-based model for the Holstein model under a deep quench, with results and benchmark analyses discussed in Section V. There we show that the trained model successfully reproduces the statistical climate of the trajectories, as evidenced by the agreement of the autocorrelation function. Finally, Section VI provides a summary and outlook. All source code and model weights are available at github.com/apning/holstein-parc.
@@ II Non-adiabatic dynamics of Holstein model
$\hat{\mathcal{H}}=-t_{\rm nn}\sum_{i}\left(\hat{c}^{{\dagger}}_{i}\hat{c}^{\,}_{i+1}+\hat{c}^{{\dagger}}_{i+1}\hat{c}^{\,}_{i}\right)-g\sum_{i}\left(\hat{n}_{i}-\frac{1}{2}\right)\hat{Q}_{i}$
@@
Here $\hat{c}^{\dagger}_{i}$ ($\hat{c}^{\,}_{i}$) denotes the creation (annihilation) operator of an electron at lattice site-$i$, $Q_{i}$ represents a local phonon degree of freedom, $P_{i}$ is the associated conjugate momentum, $m$ is an effective mass, $\Omega$ is the intrinsic oscillation frequency, and $K\equiv m\Omega^{2}$ is the force constant, finally $g$ denotes the electron-phonon coupling coefficient. The first term $\hat{\mathcal{H}}_{\mathrm{e}}$ describes the hopping of electrons between nearest neighboring sites. The second part essentially describes the Einstein phonon model, which corresponds to a set of simple harmonic oscillators each associated with a lattice site. The third term describes a local coupling between the electron density $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}^{\,}_{i}$ and the displacement of the oscillator.
The Holstein model at half-filling on various bipartite lattices exhibits a robust commensurate charge density wave (CDW) order that breaks the sublattice symmetry [44, 45, 46, 47]. In one dimension, this commensurate CDW order is characterized by a ultra-short period modulation of electron density,
$\langle\hat{n}_{i}\rangle=\overline{n}+\delta n\cos\left(\frac{\pi}{a}x_{i}\right),$ (2)
where $\langle\cdots\rangle$ denotes ground-state expectation value, $\overline{n}=1/2$ is the average density, $\delta n$ is the modulation amplitude, $a$ is the lattice constant, and $x_{i}=x_{0}+i\,a$ is the physical coordinate of site-$i$. It is worth noting that the CDW order remains robust even in the semi-classical approximation. Indeed, the semi-classical phase diagram of the CDW order obtained by a hybrid Monte Carlo method agrees very well with that obtained from determinant quantum Monte Carlo simulations [47]. Within the semi-classical approximation for the CDW dynamics, the Holstein is an example of the hybrid quantum-classical systems discussed above.
@@
[…]
@@
There are two characteristic time scales for the dynamics of the Holstein model. First, from the bandwidth of the electron tight-binding model $W=4t_{\rm nn}$, one can define a time scale $\tau_{\rm e}=\hbar/t_{\rm nn}$ for the electron dynamics. Another time scale is given by the natural frequency $\Omega$ of the local simple harmonic oscillator: $\tau_{\rm L}=1/\Omega$. The dimensionless adiabatic parameter is defined as the ratio $r=\tau_{\rm e}/\tau_{\rm L}=\hbar\Omega/t_{\rm nn}$. The electron-phonon coupling can also be characterized by a dimensionless parameter $\lambda$ as follows. First, let $Q^{*}$ be lattice distortions estimated from the balance of elastic energy and electron-phonon coupling: $KQ^{*2}\sim g\langle n\rangle Q^{*}$. Assuming electron number $\langle n\rangle\sim 1$, we obtain $Q^{*}\sim g/K$. The dimensionless parameter $\lambda=gQ^{*}/W=g^{2}/WK$ corresponds to the ratio of electron-phonon coupling to the bandwidth. For all simulations discussed below, these two dimensionless parameters are set to $r=0.4$ and $\lambda=1$. The simulation time is measured in unit of $\tau_{\rm e}$, energies are measured in units of $t_{\rm nn}$, and the lattice distortion is expressed in terms of $Q^{*}$.
@@
Figure 1: Schematic diagram of the recurrent structure for non-adiabatic dynamics of the semi-classical Holstein model. The framework is based on a single CNN for all three dynamical degrees of freedom: $Q$, $P$, and $\rho$.
Figure 2: Architecture of the standard CNN model. The components $\rho$, $Q$, and $P$ are first prepared for input. $\rho$ is split into real-valued components $\rho_{\text{real}}$ and $\rho_{\text{imag}}$. $Q$ and $P$ are both inserted into the diagonals of zero-matrices. The resulting 44-channel grid is then input into the CNN model with modified ResNet-v2 based architecture. $\rho$, $Q$, and $P$ are then extracted from the output of the CNN in the reverse manner they were input.
III Standard CNN and Shallow Quench
In this section we consider RNN models for modeling the non-adiabatic dynamics of the 1D semi-classical Holstein model triggered by a small change in the electron-lattice coupling $g$, a scenario to be called shallow quench in the following. Our exact numerical simulations find that the resultant space-time evolution is less sensitive to noises in the initial conditions. Here we present a RNN scheme based on standard CNNs for modeling such relatively deterministic dynamics. While the state of the simple harmonic oscillators on a chain is described by two 1D arrays $\{Q_{i}\}$ and $\{P_{i}\}$, the minimum dynamical variables for the many-electron system are given by a density matrix $\rho_{ij}$, due to the quantum Fermi-Dirac statistics. This means that, for the simpler case of non-interacting electrons, the 1D quantum dynamics can be effectively modeled by an effective 2D classical dynamics.
III.1 CNN-based recurrent structure
For convenience, we introduce a state vector $\mathbf{u}=\{\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3},\mathbf{w}_{4}\}$ with four separate components representing the diagonal, off-diagonal elements of the density matrix, and the position and momentum vectors of the simple harmonic oscillators, respectively. Explicitly they are defined as
$\mathbf{w}_{1}=\{\rho_{11},\rho_{22},\cdots,\rho_{LL}\},$
$\mathbf{w}_{2}=\{\rho_{12},\rho_{13},\cdots,\rho_{23},\rho_{24},\cdots,\rho_{L,L-1}\},$
$\mathbf{w}_{3}=\{Q_{1},Q_{2},\cdots,Q_{L}\},$
$\mathbf{w}_{4}=\{P_{1},P_{2},\cdots,P_{L}\}.$ (7)
We partition the density matrix $\rho_{ij}$ into diagonal ($\mathbf{w}_{1}$) and off-diagonal ($\mathbf{w}_{2}$) components to highlight their conceptual distinction: the diagonal elements $\rho_{\text{diag}}$ represent on-site electron densities, while the off-diagonal elements $\rho_{\text{off-diag}}$ encode quantum coherence.
The semi-classical dynamics of the Holstein model can be cast into a form of the standard ordinary differential equation: